OpenStudy (anonymous):

what is the approximate area of a circle with a diameter of 14 cm

OpenStudy (31356):

154cm

OpenStudy (31356):

Go to this (for more calculations) http://www.cleavebooks.co.uk/scol/calcirc.htm

OpenStudy (31356):

Medal if helped.......

OpenStudy (ybarrap):

Since they are asking for approximating, let's do without a calculator: $$ Area=\pi r^2=\pi\left (\cfrac{d}{2}\right )^2=\pi\left (\cfrac{14}{2}\right )^2=\pi~7^2\approx 3 \times 49=147~cm^2 $$ Make sense?

OpenStudy (31356):

Yea, if I were you I would do it without a calculator too :D You will learn more, from it.